OpenStudy (anonymous):

z+9 over 4=2.1 . what is z ?

OpenStudy (anonymous):



OpenStudy (anonymous):

Like that?

OpenStudy (anonymous):

z+9 = 4(2.1) Z + 9 = 8.4 z = 8.4 - 9 z = -0.6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):